OpenStudy (anonymous):

What is the solution set for |–2x + 16| – 9 = –3

OpenStudy (compassionate):

Hello, you first want to solve the equation for x. Can you do that for me, hun?

OpenStudy (anonymous):

x=11

OpenStudy (compassionate):

There you go ;)

OpenStudy (anonymous):

thats the answer to the entire thing?

OpenStudy (anonymous):

nope there is 2 answers.

OpenStudy (anonymous):

11 and -11 right?

OpenStudy (anonymous):

nope.

OpenStudy (compassionate):

No. It's actually 11 and 5. Do you know why?

OpenStudy (anonymous):

okay help me idk why

OpenStudy (anonymous):

\[\bf |-2x+16|=6 \implies -2x+16=6 \ or \ -2x+16=-6\]

OpenStudy (anonymous):

You have to cases to solve for. Since it's an absolute value, the answer could be +6 or -6 but either way it would come out as 6. So we consider both cases and solve for x in each case. We get x = 5 and x = 11.

OpenStudy (anonymous):

ohhhhh okay thank you i get it.